# A Survey on Knowledge Distillation of Large Language Models

### v3

@@ Abstract
In the era of Large Language Models (LLMs), Knowledge Distillation (KD) emerges as a pivotal methodology for transferring advanced capabilities from leading proprietary LLMs, such as GPT-4, to their open-source counterparts like LLaMA and Mistral. Additionally, as open-source LLMs flourish, KD plays a crucial role in both compressing these models, and facilitating their self-improvement by employing themselves as teachers. This paper presents a comprehensive survey of KD’s role within the realm of LLM, highlighting its critical function in imparting advanced knowledge to smaller models and its utility in model compression and self-improvement. Our survey is meticulously structured around three foundational pillars: algorithm, skill, and verticalization – providing a comprehensive examination of KD mechanisms, the enhancement of specific cognitive abilities, and their practical implications across diverse fields. Crucially, the survey navigates the intricate interplay between data augmentation (DA) and KD, illustrating how DA emerges as a powerful paradigm within the KD framework to bolster LLMs’ performance. By leveraging DA to generate context-rich, skill-specific training data, KD transcends traditional boundaries, enabling open-source models to approximate the contextual adeptness, ethical alignment, and deep semantic insights characteristic of their proprietary counterparts. This work aims to provide an insightful guide for researchers and practitioners, offering a detailed overview of current methodologies in knowledge distillation and proposing future research directions. By bridging the gap between proprietary and open-source LLMs, this survey underscores the potential for more accessible, efficient, and powerful AI solutions. Most importantly, we firmly advocate for compliance with the legal terms that regulate the use of LLMs, ensuring ethical and lawful application of KD of LLMs. An associated Github repository is available at https://github.com/Tebmer/Awesome-Knowledge-Distillation-of-LLMs.
@@ 1 Introduction
In the evolving landscape of artificial intelligence (AI), proprietary [^1] For simplicity, we use ‘proprietary’ to represent both versatile yet close-source LLMs like GPT-4 and open-source yet huge LLMs like LLaMA-2-70B, which encapsulate rich knowledge with a large number of parameters. Large Language Models (LLMs) such as GPT-3.5 Ouyang et al. (2022), GPT-4 OpenAI et al. (2023), Gemini Team et al. (2023) and Claude [^2] https://www.anthropic.com/claude-in-slack have emerged as groundbreaking technologies, reshaping our understanding of natural language processing (NLP). These models, characterized by their vast scale and complexity, have unlocked new realms of possibility, from generating human-like text to offering sophisticated problem-solving capabilities. The core significance of these LLMs lies in their emergent abilities Wei et al. (2022a, b); Xu et al. (2024a), a phenomenon where the models display capabilities beyond their explicit training objectives, enabling them to tackle a diverse array of tasks with remarkable proficiency. Their deep understanding of context, nuance, and the intricacies of human language enables them to excel in a wide array of applications, from creative content generation to complex problem-solving OpenAI et al. (2023); Liang et al. (2022). The potential of these models extends far beyond current applications, promising to revolutionize industries, augment human creativity, and redefine our interaction with technology.
@@
The escalating need for a comprehensive survey on the knowledge distillation of LLMs stems from the rapidly evolving landscape of AI OpenAI et al. (2023); Team et al. (2023) and the increasing complexity of these models. As AI continues to penetrate various sectors, the ability to efficiently and effectively distill knowledge from proprietary LLMs to open-source ones becomes not just a technical aspiration but a practical necessity. This need is driven by the growing demand for more accessible, cost-effective, and adaptable AI solutions that can cater to a diverse range of applications and users. A survey in this field is vital for synthesizing the current methodologies, challenges, and breakthroughs in knowledge distillation. It may serve as a beacon for researchers and practitioners alike, guiding them through the intricate process of distilling complex AI capabilities into more manageable and accessible forms. Moreover, such a survey can illuminate the path forward, identifying gaps in current techniques and proposing directions for future research.
Figure 2: An overview of this survey on knowledge distillation of large language models. Note that ‘Section’ is abbreviated as ‘Sec.’ in this figure. $\text{RM}_{S}(\cdot)$ denotes the student reward model.
@@
Lion Jiang et al. (2023b), PERsD Chen et al. (2023a), GKD Agarwal et al. (2024), leaf, text width=31em ] ] [ Self-Knowledge [ Self-Instruct Wang et al. (2022a), Self-Align Sun et al. (2024b), RLCD Yang et al. (2024a),
@@
Self-Align Sun et al. (2024b), UltraFeedback Cui et al. (2023a), RLCD Yang et al. (2024a), leaf, text width=31em ] ] ] [ Agent [ Tool Using [ Toolformer Schick et al. (2023), Graph-ToolFormer Zhang (2023), Gorilla Patil et al. (2023),
@@ 2.1 Comparing Traditional Recipe
In contrast, the advent of LLMs has revolutionized the knowledge distillation landscape. The current era of knowledge distillation in LLMs shifts the focus from mere architecture compression to the more nuanced process of knowledge elicitation and transfer Taori et al. (2023); Chaudhary (2023); Tunstall et al. (2023). This paradigm change is largely due to the expansive and deep-seated knowledge that LLMs like GPT-4 and Gemini possess. And the inaccessible parameters of LLMs make it hard to compress them by using pruning Han et al. (2016) or quantization Liu et al. (2023a) techniques. Unlike the earlier era, where the goal was to replicate the output behavior of the teacher model or reduce the model size , the current focus in LLM-based knowledge distillation is to extract and transfer the rich, nuanced understanding that these models have developed.
@@ 2.2 Relation to Data Augmentation (DA)
In the era of LLMs, Data Augmentation (DA) Wang et al. (2022a); Ye et al. (2022) emerges as a critical paradigm integral to the process of knowledge distillation. Unlike traditional DA techniques such as paraphrasing Gangal et al. (2022) or back-translation Longpre et al. (2019), which primarily aim at expanding the training dataset in a somewhat mechanical manner. DA within the context of LLMs focuses on the generation of novel, context-rich training data tailored to specific domains and skills. This innovation is driven by the unique capabilities of LLMs to generate coherent, diverse, and intricate data samples that closely mimic the nuanced understanding and cognitive abilities of human experts in various fields.
@@
The significance and necessity of DA for achieving KD in the LLM era cannot be overstated. DA acts as a force multiplier, enabling the distilled models to acquire and refine capabilities that would otherwise require exponentially larger datasets and computational resources. It facilitates a more nuanced and effective transfer of knowledge, focusing on the qualitative aspects of learning rather than quantitative expansion. This strategic use of DA within KD processes underscores a pivotal shift towards a more efficient, sustainable, and accessible approach to harnessing the power of LLMs. It empowers open-source models with the ability to approximate the contextual adeptness, ethical alignment, and deep semantic insights characteristic of their proprietary counterparts, thereby democratizing access to advanced AI capabilities and fostering innovation across a broader spectrum of applications and users.
@@ 2.3 Survey Scope
KD Algorithms. This segment focuses on the technical foundations and methodologies of knowledge distillation. It includes an in-depth exploration of the processes involved in constructing knowledge from teacher models (e.g., proprietary LLMs) and integrating this knowledge into student models (e.g., open-source LLMs). Under the umbrella of ‘knowledge’, we delve into strategies such as labeling Hsieh et al. (2023), expansion Taori et al. (2023), curation Gunasekar et al. (2023), feature understanding Agarwal et al. (2024), feedback mechanisms Tunstall et al. (2023), and self-knowledge generation Wang et al. (2022a). This exploration seeks to uncover the various ways in which knowledge can be identified, expanded, and curated for effective distillation. The ‘distillation’ subsection examines learning approaches like supervised fine-tuning (SFT) Wang et al. (2022a), divergence minimization Agarwal et al. (2024), reinforcement learning techniques Cui et al. (2023a), and rank optimization strategies Tunstall et al. (2023). This analysis aims to illuminate how these algorithms facilitate the transfer of knowledge, ensuring that open-source models can replicate and, in some cases, surpass the capabilities of their proprietary counterparts.
@@
Skill Distillation. This facet examines the specific competencies and capabilities enhanced through KD. It encompasses detailed discussions on context following Taori et al. (2023); Luo et al. (2023c), with subtopics like instruction following and retrieval-augmented generation (RAG) Capability. In the realm of alignment Mitra et al. (2023); Tunstall et al. (2023), the survey investigates thinking patterns, persona/preference modeling, and value alignment. The ‘agent’ category delves into skills such as Tool Using and Planning. NLP task specialization Dai et al. (2023a); Jung et al. (2023); Chaudhary (2023) is scrutinized through lenses like natural language understanding (NLU), natural language generation (NLG), information retrieval, recommendation systems, text generation evaluation, and code generation. Finally, the survey addresses multi-modality Liu et al. (2023e); Zhao et al. (2023b), exploring how KD enhances LLMs’ ability to interpret and integrate multiple forms of input, enriching their utility and applicability across various contexts.
@@
Verticalization Distillation. This section assesses the application of KD across diverse vertical domains, offering insights into how distilled LLMs can be tailored for specialized fields such as Law LAW (2023), Medical & Healthcare Wang et al. (2023a), Finance Zhang and Yang (2023), Science Zhang et al. (2024), among others. This exploration not only showcases the practical implications of KD techniques but also highlights their transformative impact on domain-specific AI solutions. Through detailed analysis and examples, this part aims to demonstrate the versatility and efficacy of KD in adapting LLMs to meet the nuanced demands of different industries, thus contributing to the broader AI and ML ecosystem.
@@
By navigating through these facets, this survey endeavors to provide an extensive and nuanced analysis of knowledge distillation in the era of LLMs. It serves as a guide for researchers, practitioners, and enthusiasts in the field, shedding light on current methodologies, challenges, and opportunities for innovation in this rapidly evolving domain.
@@ 2.4 Distillation Pipeline in LLM Era
I. Target Skill or Domain Steering Teacher LLM. The first stage involves directing the teacher LLM towards a specific target skill or domain. This is achieved through carefully crafted instructions or templates that guide the LLM’s focus. These instructions are designed to elicit responses that demonstrate the LLM’s proficiency in a particular area, be it a specialized domain like healthcare or law, or a skill such as reasoning or language understanding. The objective here is to utilize the teacher LLM’s extensive training and nuanced capabilities to generate outputs that are rich in the specific knowledge or skills desired for the student model.
@@ 3.1.4 Feature
In summary, feature knowledge offers a more transparent alternative to black-box methods, allowing for deeper insight into and control over the distillation process. By utilizing feature knowledge from teacher LLMs, such as output distributions and intermediate layer features, white-box approaches enable a more nuanced transfer of information. While showing promise, especially in smaller models, its application is not suitable for black-box LLMs where internal parameters are inaccessible. Furthermore, student models distilled from white-box LLMs may underperform compared to their black-box counterparts, as the black-box teacher LLMs (e.g. GPT-4) tend to be more powerful.
@@ 3.1.6 Self-Knowledge
where $\phi_{\text{sk}}(\cdot)$ is a generalized function that represents an additional process to the self-generated outputs $y$, which could include but is not limited to filtering, rewarding, or any other mechanisms for enhancing or evaluating $y$. It could be governed by external tools or the student itself $\theta_{S}$. Recent research in this area has proposed various innovative methodologies to elicit self-knowledge, demonstrating its potential for creating more efficient and autonomous learning systems. Allen-Zhu and Li (2020); Wang et al. (2022a); Sun et al. (2024b); Yang et al. (2024a); Jung et al. (2023); Huang et al. (2023a); Gulcehre et al. (2023); Yuan et al. (2024a); Xu et al. (2023b); Zelikman et al. (2022); Chen et al. (2024a); Zheng et al. (2024); Li et al. (2024c); Zhao et al. (2024); Singh et al. (2023); Chen et al. (2024c); Hosseini et al. (2024)
A notable example of this methodology is Self-Instruct Wang et al. (2022a), which utilizes GPT-3 for data augmentation through the Expansion approach, generating additional data samples to enhance the dataset. This enriched dataset subsequently fine-tunes the original model. Other methods aim to elicit targeted knowledge from student models by modifying prompts, and leveraging these data for further refinement. In Self-Align Sun et al. (2024b), they find that models fine-tuned by Self-Instruct data tend to generate short or indirect responses. They prompt this model with verbose instruction to produce in-depth and detailed responses. Then, they employ context-distillation Askell et al. (2021) to distill these responses paired with non-verbose instructions back to the model. Similarly, RLCD Yang et al. (2024a) introduces the use of contrasting prompts to generate preference pairs from an unaligned LLM, encompassing both superior and inferior examples. A preference model trained on these pairs then guides the enhancement of the unaligned model through reinforcement learning. Several other approaches employ filtering methods to refine self-generated data. For example, Impossible Distillation Jung et al. (2023) targets sentence summarization tasks, implementing filters based on entailment, length, and diversity to screen self-generated summaries. LMSI Huang et al. (2023a) generates multiple CoT reasoning paths and answers for each question, and then retains only those paths that lead to the most consistent answer.
@@ 3.2.2 Divergence and Similarity
The specific form of $D$ varies depending on the type of divergence employed. Table I outlines the functional forms of $D$ for different divergence measures. The commonly-used standard KD objectives essentially minimize the approximated forward Kullback-Leibler divergence (KLD) between the teacher and the student distribution Sanh et al. (2019); Wen et al. (2023); Timiryasov and Tastet (2023); Liang et al. (2023a); Chen et al. (2024d) , which forces $p_{S}$ to cover all the modes of $p_{T}$. However, when a student model is unable to learn all modes of a highly complex teacher, the resultant “mode-covering” behavior might cause the student to assign probability mass to tokens with low probability under the teacher’s distribution (cf. Figure 6 blue curve). This mode-covering phenomenon can potentially lead to hallucinations and low-quality generations. Alternatively, mode-seeking divergences like reverse KL prioritize tokens where the teacher assigns high probabilities (cf. Figure 6 green curve). This approach can mitigate the risk of low-quality outputs, fostering more accurate generations. However, it often does so at the cost of reduced diversity. Gu et al. (2024) adopt reverse KL divergence to prevent students from overestimating low-probability regions of the teacher’s distribution, employing Policy Gradient methods for optimization. Both Agarwal et al. (2024) and Sason and Verdú (2016) assess the efficacy of different divergence functions in LLM distillation, finding the optimal divergence to be task-dependent. For instance, forward KL divergence is more suitable for tasks like Machine Translation, where the output has fewer modes or variations, while reverse KL divergence is preferable for tasks like dialogue generation and instruction tuning, which involve multiple modes and a wider range of potential responses. Thus, the nature of the task significantly influences the selection of the divergence function for optimal performance.

### v4

@@ Abstract
In the era of Large Language Models (LLMs), Knowledge Distillation (KD) emerges as a pivotal methodology for transferring advanced capabilities from leading proprietary LLMs, such as GPT-4, to their open-source counterparts like LLaMA and Mistral. Additionally, as open-source LLMs flourish, KD plays a crucial role in both compressing these models, and facilitating their self-improvement by employing themselves as teachers. This paper presents a comprehensive survey of KD’s role within the realm of LLM, highlighting its critical function in imparting advanced knowledge to smaller models and its utility in model compression and self-improvement. Our survey is meticulously structured around three foundational pillars: algorithm, skill, and verticalization – providing a comprehensive examination of KD mechanisms, the enhancement of specific cognitive abilities, and their practical implications across diverse fields. Crucially, the survey navigates the interaction between data augmentation (DA) and KD, illustrating how DA emerges as a powerful paradigm within the KD framework to bolster LLMs’ performance. By leveraging DA to generate context-rich, skill-specific training data, KD transcends traditional boundaries, enabling open-source models to approximate the contextual adeptness, ethical alignment, and deep semantic insights characteristic of their proprietary counterparts. This work aims to provide an insightful guide for researchers and practitioners, offering a detailed overview of current methodologies in knowledge distillation and proposing future research directions. By bridging the gap between proprietary and open-source LLMs, this survey underscores the potential for more accessible, efficient, and powerful AI solutions. Most importantly, we firmly advocate for compliance with the legal terms that regulate the use of LLMs, ensuring ethical and lawful application of KD of LLMs. An associated Github repository is available at https://github.com/Tebmer/Awesome-Knowledge-Distillation-of-LLMs.
@@ 1 Introduction
In the evolving landscape of artificial intelligence (AI), proprietary [^1] For simplicity, we use ‘proprietary’ to represent both versatile yet close-source LLMs like GPT-4 and open-source yet huge LLMs like LLaMA-2-70B, which encapsulate rich knowledge with a large number of parameters. Large Language Models (LLMs) such as GPT-3.5 Ouyang et al. (2022), GPT-4 OpenAI et al. (2023), Gemini Team et al. (2023) and Claude [^2] https://www.anthropic.com/claude-in-slack have emerged as groundbreaking technologies, reshaping our understanding of natural language processing (NLP). These models, characterized by their vast scale and complexity, have unlocked new realms of possibility, from generating human-like text to offering sophisticated problem-solving capabilities. The core significance of these LLMs lies in their emergent abilities Wei et al. (2022a, b); Xu et al. (2024a), a phenomenon where the models display capabilities beyond their explicit training objectives, enabling them to tackle a diverse array of tasks with remarkable proficiency. These models excel in understanding and generation, driving applications from creative generation to complex problem-solving OpenAI et al. (2023); Liang et al. (2022). The potential of these models extends far beyond current applications, promising to revolutionize industries, augment human creativity, and redefine our interaction with technology.
@@
The escalating need for a comprehensive survey on the knowledge distillation of LLMs stems from the rapidly evolving landscape of AI OpenAI et al. (2023); Team et al. (2023) and the increasing complexity of these models. As AI continues to penetrate various sectors, the ability to efficiently and effectively distill knowledge from proprietary LLMs to open-source ones becomes not just a technical aspiration but a practical necessity. This need is driven by the growing demand for more accessible, cost-effective, and adaptable AI solutions that can cater to a diverse range of applications and users. A survey in this field is vital for synthesizing the current methodologies, challenges, and breakthroughs in knowledge distillation. It may serve as a beacon for researchers and practitioners alike, guiding them to distill complex AI capabilities into more manageable and accessible forms. Moreover, such a survey can illuminate the path forward, identifying gaps in current techniques and proposing directions for future research.
[…]
@@
Lion Jiang et al. (2023b), PERsD Chen et al. (2023a), GKD Agarwal et al. (2024), leaf, text width=31em ] ] [ Self-Knowledge [ Self-Instruct Wang et al. (2022a), Self-Align Sun et al. (2024b), RLCD Yang et al. (2024),
@@
Self-Align Sun et al. (2024b), UltraFeedback Cui et al. (2023a), RLCD Yang et al. (2024), leaf, text width=31em ] ] ] [ Agent [ Tool Using [ Toolformer Schick et al. (2023), Graph-ToolFormer Zhang (2023), Gorilla Patil et al. (2023),
@@ 2.1 Comparing Traditional Recipe
In contrast, the advent of LLMs has revolutionized the knowledge distillation landscape. The current era of knowledge distillation in LLMs shifts the focus from mere architecture compression to knowledge elicitation and transfer Taori et al. (2023); Chaudhary (2023); Tunstall et al. (2023). This paradigm change is largely due to the expansive and deep-seated knowledge that LLMs like GPT-4 and Gemini possess. And the inaccessible parameters of LLMs make it hard to compress them by using pruning Han et al. (2016) or quantization Liu et al. (2023a) techniques. Unlike the earlier era, where the goal was to replicate the output behavior of the teacher model or reduce the model size, the current focus in LLM-based knowledge distillation is to elicit the specific knowledge these models have.
@@ 2.2 Relation to Data Augmentation (DA)
In the era of LLMs, Data Augmentation (DA) Wang et al. (2022a); Ye et al. (2022) emerges as a critical paradigm integral to the process of knowledge distillation. Unlike traditional DA techniques such as paraphrasing Gangal et al. (2022) or back-translation Longpre et al. (2019), which primarily aim at expanding the training dataset in a somewhat mechanical manner, DA within the context of LLMs focuses on the generation of novel, context-rich training data tailored to specific domains and skills.
@@
DA acts as a force multiplier, enabling the distilled models to acquire and refine capabilities that would otherwise require exponentially larger datasets and computational resources. It facilitates a more effective transfer of knowledge, focusing on the qualitative aspects of learning rather than quantitative expansion. This strategic use of DA within KD processes underscores a pivotal shift towards a more efficient, sustainable, and accessible approach to harnessing the power of LLMs. It empowers open-source models with the ability to approximate the contextual adeptness, ethical alignment, and deep semantic insights characteristic of their proprietary counterparts, thereby democratizing access to advanced AI capabilities and fostering innovation across a broader spectrum of applications and users.
@@ 2.3 Survey Scope
KD Algorithms. This segment focuses on the technical foundations and methodologies of knowledge distillation. It includes an in-depth exploration of the processes involved in constructing knowledge from teacher models (e.g., proprietary LLMs) and integrating this knowledge into student models (e.g., open-source LLMs). Under the umbrella of ‘knowledge’, we delve into strategies such as labeling Hsieh et al. (2023), expansion Taori et al. (2023), curation Gunasekar et al. (2023), feature understanding Agarwal et al. (2024), feedback mechanisms Tunstall et al. (2023), and self-knowledge generation Wang et al. (2022a). This exploration seeks to uncover the various ways in which knowledge can be identified, expanded, and curated for effective distillation. The ‘distillation’ subsection examines learning approaches like supervised fine-tuning (SFT) Wang et al. (2022a), divergence minimization Agarwal et al. (2024), reinforcement learning techniques Cui et al. (2023a), and rank optimization strategies Tunstall et al. (2023). Together, these techniques demonstrate how KD enables open-source models to obtain knowledge from proprietary ones.
@@
Skill Distillation. This facet examines the specific competencies and capabilities enhanced through KD. It encompasses detailed discussions on context following Taori et al. (2023); Luo et al. (2023c), with subtopics like instruction following and retrieval-augmented generation (RAG) Capability. In the realm of alignment Mitra et al. (2023); Tunstall et al. (2023), the survey investigates thinking patterns, persona/preference modeling, and value alignment. The ‘agent’ category delves into skills such as Tool Using and Planning. NLP task specialization Dai et al. (2023a); Jung et al. (2023); Chaudhary (2023) is scrutinized through lenses like natural language understanding (NLU), natural language generation (NLG), information retrieval, recommendation systems, text generation evaluation, and code generation. Finally, the survey addresses multi-modality Liu et al. (2023e); Zhao et al. (2023b), exploring how KD enhances LLMs’ ability to integrate multiple forms of input.
@@
Verticalization Distillation. This section assesses the application of KD across diverse vertical domains, offering insights into how distilled LLMs can be tailored for specialized fields such as Law LAW (2023), Medical & Healthcare Wang et al. (2023a), Finance Zhang and Yang (2023), Science Zhang et al. (2024), among others. This exploration not only showcases the practical implications of KD techniques but also highlights their transformative impact on domain-specific AI solutions.
@@
Through these facets, this survey provides a comprehensive analysis of KD in LLMs, guiding researchers and practitioners through methodologies, challenges, and opportunities in this rapidly evolving domain.
@@ 2.4 Distillation Pipeline in LLM Era
I. Target Skill or Domain Steering Teacher LLM. The first stage involves directing the teacher LLM towards a specific target skill or domain. This is achieved through carefully crafted instructions or templates that guide the LLM’s focus. These instructions are designed to elicit responses that demonstrate the LLM’s proficiency in a particular area, be it a specialized domain like healthcare or law, or a skill such as reasoning or language understanding.
@@ 3.1.4 Feature
In summary, feature knowledge offers a more transparent alternative to black-box methods, allowing for deeper insight into and control over the distillation process. By utilizing feature knowledge from teacher LLMs, such as output distributions and intermediate layer features, white-box approaches enable richer knowledge transfer. While showing promise, especially in smaller models, its application is not suitable for black-box LLMs where internal parameters are inaccessible. Furthermore, student models distilled from white-box LLMs may underperform compared to their black-box counterparts, as the black-box teacher LLMs (e.g. GPT-4) tend to be more powerful.
@@ 3.1.6 Self-Knowledge
where $\phi_{\text{sk}}(\cdot)$ is a generalized function that represents an additional process to the self-generated outputs $y$, which could include but is not limited to filtering, rewarding, or any other mechanisms for enhancing or evaluating $y$. It could be governed by external tools or the student itself $\theta_{S}$. Recent research in this area has proposed various innovative methodologies to elicit self-knowledge, demonstrating its potential for creating more efficient and autonomous learning systems. Allen-Zhu and Li (2020); Wang et al. (2022a); Sun et al. (2024b); Yang et al. (2024); Jung et al. (2023); Huang et al. (2023a); Gulcehre et al. (2023); Yuan et al. (2024a); Xu et al. (2023b); Zelikman et al. (2022); Chen et al. (2024a); Zheng et al. (2024); Li et al. (2024c); Zhao et al. (2024); Singh et al. (2023); Chen et al. (2024c); Hosseini et al. (2024)
A notable example of this methodology is Self-Instruct Wang et al. (2022a), which utilizes GPT-3 for data augmentation through the Expansion approach, generating additional data samples to enhance the dataset. This enriched dataset subsequently fine-tunes the original model. Other methods aim to elicit targeted knowledge from student models by modifying prompts, and leveraging these data for further refinement. In Self-Align Sun et al. (2024b), they find that models fine-tuned by Self-Instruct data tend to generate short or indirect responses. They prompt this model with verbose instruction to produce in-depth and detailed responses. Then, they employ context-distillation Askell et al. (2021) to distill these responses paired with non-verbose instructions back to the model. Similarly, RLCD Yang et al. (2024) introduces the use of contrasting prompts to generate preference pairs from an unaligned LLM, encompassing both superior and inferior examples. A preference model trained on these pairs then guides the enhancement of the unaligned model through reinforcement learning. Several other approaches employ filtering methods to refine self-generated data. For example, Impossible Distillation Jung et al. (2023) targets sentence summarization tasks, implementing filters based on entailment, length, and diversity to screen self-generated summaries. LMSI Huang et al. (2023a) generates multiple CoT reasoning paths and answers for each question, and then retains only those paths that lead to the most consistent answer.
@@ 3.2.2 Divergence and Similarity
The specific form of $D$ varies depending on the type of divergence employed. Table I outlines the functional forms of $D$ for different divergence measures. The commonly-used standard KD objectives essentially minimize the approximated forward Kullback-Leibler divergence (KLD) between the teacher and the student distribution Sanh et al. (2019); Wen et al. (2023); Timiryasov and Tastet (2023); Liang et al. (2023a); Chen et al. (2024d) , which forces $p_{S}$ to cover all the modes of $p_{T}$. However, when a student model is unable to learn all modes of a highly complex teacher, the resultant “mode-covering” behavior might cause the student to assign probability mass to tokens with low probability under the teacher’s distribution (cf. Figure 6 blue curve). This mode-covering phenomenon can potentially lead to hallucinations and low-quality generations. Alternatively, mode-seeking divergences like reverse KL prioritize tokens where the teacher assigns high probabilities (cf. Figure 6 green curve). This approach can mitigate the risk of low-quality outputs, fostering more accurate generations. However, it often does so at the cost of reduced diversity. Gu et al. (2024) adopt reverse KL divergence to prevent students from overestimating low-probability regions of the teacher’s distribution, employing Policy Gradient methods for optimization. Both Agarwal et al. (2024) and Sason and Verdú (2016) assess the effect of different divergence functions in LLM distillation, finding the optimal divergence to be task-dependent. For instance, forward KL divergence is more suitable for tasks like Machine Translation, where the output has fewer modes or variations, while reverse KL divergence is preferable for tasks like dialogue generation and instruction tuning, which involve multiple modes and a wider range of potential responses. Thus, the nature of the task significantly influences the selection of the divergence function for optimal performance.